Calculate the percent yield if 15.0 g of nitrogen gas is produced from the decomposition of 23.6 g of sodium azide (NaN3) according to the following chemical equation. Give your answer as a %.

2 NaN3(s) → 3 N2(g) + 2 Na(s)
Answer #1

Molar mass of NaN3,

MM = 1*MM(Na) + 3*MM(N)

= 1*22.99 + 3*14.01

= 65.02 g/mol

mass of NaN3 = 23.6 g

mol of NaN3 = (mass)/(molar mass)

= 23.6/65.02

= 0.363 mol

Balanced chemical equation is:

2 NaN3 -> 2 Na + 3 N2

According to balanced equation

mol of N2 formed = (3/2)* moles of NaN3

= (3/2)*0.363

= 0.5444 mol

Molar mass of N2 = 28.02 g/mol

mass of N2 = number of mol * molar mass

= 0.5444*28.02

= 15.26 g

% yield = actual mass*100/theoretical mass

= 15*100/15.26

= 98.33 %

Answer: 98.3 %